OpenStudy (anonymous):

y=6-x and y=x-2 a (2,4) b (-4,2) c (4,2) d no solution

OpenStudy (anonymous):

y=6-x rearrange this equation to look like -6=-x-y Plug the second equation into the first equation to look like this -6=-x-(x-2) Simplify to look like this -6=-2x + 2 -6-2=-2x -8=-2x 4=x Plug in x=4 into the second equation and you get this y= 4-2 y=2 So your answer is c